OpenStudy (anonymous):

Decide whether each equation has one solution, no solutions, or infinitely many solutions. 1. 2(x - 3) = 2x A. one solution B. no solutions C. infinitely many solutions 2. 3(y - 3) = 2y - 9 + y A. one solution B. no solutions C. infinitely many solutions 3. 10x - 2 - 6x = 3x - 2 + x A. one solution B. no solutions C. infinitely many solutions 4. 4(x + 3) + 2x = x - 8 A. one solution B. no solutions C. infinitely many solutions

OpenStudy (anonymous):

1) No solution because 2x-6≠2x 2) Infinite solutions because 3y-9= 3y-9 3) Infinite solutions because 4x- 2= 4x-2 4) One solution

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

1. 2(x - 3) = 2x -- distribute through the parenthesis 2x - 6 = 2x -- add 6 to both sides 2x = 2x + 6 -- subtract 2x from both sides 2x - 2x = 6 0 = 6 (incorrect) If you answer comes out not equal, then there is no solution 2. 3(y - 3) = 2y - 9 + y -- distribute and combine 3y - 9 = 3y - 9 -- subtract 3y 3y - 3y - 9 = - 9 -- add 9 3y - 3y = -9 + 9 0 = 0 (correct) If you answer comes out equal, then there is infinite solutions 3. 10x - 2 - 6x = 3x - 2 + x -- combine like terms 4x - 2 = 4x - 2 same thing here...it is gonna come out to 0 = 0 making it have infinite solutions 4. 4(x + 3) + 2x = x - 8 4x + 12 + 2x = x - 8 6x + 12 = x - 8 6x - x = -8 - 12 5x = - 20 x = -4 ONE SOLUTION

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

heyy @sassydiva321

OpenStudy (anonymous):

This helped with one of my assignments that I just had LOL

OpenStudy (anonymous):

hahaha glad it could help cx @sassydiva321

OpenStudy (alexis55):

which one was correct